# Cycles in spherical Deligne complexes and application to $K(\pi,1)$-conjecture for Artin groups

### v1

@@ 1.2 Main results
(=Theorem 11.5) A Dynkin diagram $\Lambda$ is called an extended star if
  1. 1.
    $\Lambda$ is a tree with a vertex $s\in\Lambda$ such that each edge containing $s$ has label $=3$;
  2. 2.
    each component of $\Lambda\setminus\{s\}$ either contains only one vertex, or contains only one edge whose label is $4$.
If $\Lambda$ is an extended star, then $A_{\Lambda}$ satisfies the $K(\pi,1)$-conjecture.
Another type of result concerns Artin groups with fewer generators. Artin groups with $\leq 3$ generators were understood [Del72, CD95a, DPS22]. However, Artin groups with $4$ generators already contain many examples whose $K(\pi,1)$ were not understood before.
@@
(=Corollary 9.3) Let $A_{S}$ be an Artin group with $\leq 4$ generators. Suppose the Dynkin diagram of $A_{S}$ is not $[3,5,3]$. Then $A_{S}$ satisfies the $K(\pi,1)$-conjecture.
One can obtain more general new examples of high dimensional Artin groups satisfying $K(\pi,1)$ by combining the results in this article, and [Hua23]. Here is a simple example.
@@
(=Corollary 10.9) Given an Artin group $A_{S}$ with a complete presentation graph. If its Dynkin diagram $\Lambda$ is a tree Dynkin diagram with a collection of open edges $E$ with label $\geq 6$ such that each component of $\Lambda\setminus(\cup_{e\in E}\{e\})$ is either spherical or a star $K_{1,n}$. Then $A_{\Lambda}$ satisfies the $K(\pi,1)$-conjecture.
@@
Let $\Lambda$ be a connected Dynkin diagram satisfying the assumptions in one of the previous theorems. Then $A_{\Lambda}$ is torsion free and it has a trivial center as long as $\Lambda$ is not spherical.
We end this section by making a summary of all the examples of hyperbolic type Artin groups with known $K(\pi,1)$-conjecture which are either Lannér or quasi-Lannér, combining results from [CD95a, Cha04, DPS22, HH23, Hua23] as well as this article.
@@
Suppose $\Lambda$ is a Lannér or quasi-Lannér Dynkin diagram of hyperbolic type. Then $K(\pi,1)$-conjecture is known for $A_{\Lambda}$ when either $\Lambda$ is associated with a reflection group acting on $\mathbb{H}^{n}$ with $n\leq 3$ except $\Lambda=[3,5,3]$, or $\Lambda$ belongs to one of the following diagrams in Figure 3.
Figure 3: Diagram with dimension $\geq 4$ with known $K(\pi,1)$.
1.3 Comment on the proofs
We will first review the strategy in [Hua23], then explain what is new in this article.
Brief recap from the [Hua23]
One general principle to prove the $K(\pi,1)$-conjecture for Artin group $A_{S}$ inductively, proposed in [CD95a], is to construct a suitable contractible complex $X$ where $A_{S}$ acts with cell-stabilizers being smaller Artin groups whose $K(\pi,1)$-conjecture are known. There are several different possible complexes one can consider. Recall that the Artin complex, introduced in [CD95a] and further studied in [GP12b, CMV20], is defined as follows. For each $s\in S$, let $A_{\hat{s}}$ be the standard parabolic subgroup generated by $S\setminus\{s\}$. Let $\Delta_{S}$ be the simplicial complex whose vertex set corresponds to left cosets of $\{A_{\hat{s}}\}_{s\in S}$. Moreover, a collection of vertices spans a simplex if the associated cosets have a nonempty common intersection. The Artin complex is an analogue of Coxeter complex in the setting of Artin group. The definition of a Coxeter complex $\mathsf{C}_{S}$ of a Coxeter group $W_{S}$ is almost identical to the Artin complex, except one replaces $A_{\hat{s}}$ by $W_{\hat{s}}$, which is the standard parabolic subgroup of $W_{S}$ generated by $S\setminus\{s\}$. Each vertex of $\mathsf{C}_{S}$ or $\Delta_{S}$ corresponding to a left coset of $W_{\hat{s}}$ or $A_{\hat{s}}$ has a type, which is defined to be $\hat{s}=S\setminus\{s\}$. The following gives an inductive way to prove the $K(\pi,1)$-conjecture.

### v3

@@ 1.3 Application to the $K(\pi,1)$-conjecture
(=Corollary 10.4) Let $W_{S}$ be a Coxeter group acting properly on $\mathbb{H}^{3}$ by isometries with finite covolume. Assume $W_{S}\neq[3,5,3]$. Then the $K(\pi,1)$-conjecture holds for the associated Artin group $A_{S}$.
Here $[3,5,3]$ denotes the Coxeter group whose Coxeter diagram is a linear graph with three edges, such that consecutive edges are labeled $3,5,3$. This remaining group has some exceptional feature and will be treated separately in [HP25] along with other classes. Section 1.2 also implies new results in dimension 4, with some restriction on the shape of the fundamental domain.
@@
(=Corollary 12.5) Let $W_{S}$ be a Coxeter group acting properly on $\mathbb{H}^{n}$ by isometries with its fundamental domain being a finite volume non-compact simplex. Suppose $n\leq 4$. Then the $K(\pi,1)$-conjecture holds for the associated Artin group $A_{S}$.
For higher dimensional Artin groups, we prove the following.
Theorem 1.8.
(=Theorem 11.12) Let $A_{S}$ be an Artin group such that its Coxeter diagram is complete bipartite. Then $A_{S}$ satisfies the $K(\pi,1)$-conjecture.
A complete bipartite graph $\Lambda$ is a graph whose vertices can be split into two non-empty subsets $V_{1}$ and $V_{2}$ so that $\Lambda$ has an edge between two vertices if and only if one of these vertices belongs to $V_{1}$ and the other belongs to $V_{2}$. We write $K_{m,n}$ to denote the complete bipartite graph with $m$ vertices on one subset and $n$ vertices in the other.
In each dimension $\geq 5$, there are infinitely many free-of-infinity Artin groups satisfying the assumption of Theorem 1.8 whose $K(\pi,1)$-conjecture were not previously known. The edge labels of the diagrams in Theorem 1.8 is allowed to be arbitrary. While a few instances of Artin groups in Theorem 1.8 are of hyperbolic type, most of them are not, as hyperbolic type Artin groups exist only up to a certain dimension.
Artin groups with complete bipartite Coxeter diagrams were used by McCammond [McC17, Section 8] to demonstrate how little we understand about Artin groups. When the Coxeter diagram is the star $K_{1,n}$ and all edges are labeled by $3$, the only previously known cases of $K(\pi,1)$-conjecture were $K_{1,1}$ (type $A_{2}$) and $K_{1,2}$ (type $A_{3}$) in [FN62], $K_{1,3}$ (type $D_{4}$) in [Bri06, Del72] and $K_{1,4}$ (type $\widetilde{D}_{4}$) recently in [PS21]. The $K(\pi,1)$-conjecture for the next diagram $K_{1,5}$, whose Coxeter group acts on $\mathbb{H}^{5}$ with finite volume fundamental domain, was already not known before.
We now remark briefly on the passage from results in Section 1.2 to $K(\pi,1)$-results. An Artin group $A_{S}$ is almost spherical, if it is not spherical, but $A_{S\setminus\{s\}}$ is spherical for any $s\in S$. Coxeter diagrams of irreducible almost spherical Artin groups are either contained in the last two columns of Figure 1, or in [Dav12, Table 6.2]. A sub-diagram $\Lambda^{\prime}$ of a Coxeter diagram $\Lambda$ is almost spherical, if $A_{\Lambda^{\prime}}$ is almost spherical.
The approach in [Hua24a] studies each Artin group via the information of what are the almost spherical sub-diagrams of its Coxeter diagram, and how these sub-diagrams sit inside the ambient diagram. This information can be further related to the study of short cycles in spherical Deligne complexes. Hence combining [Hua24a] and Section 1.2 output results on $K(\pi,1)$-conjecture for Artin groups such that the types and relative positions of almost spherical sub-diagrams in their Coxeter diagrams satisfy certain restrictions, depending on the list of short cycles in spherical Deligne complexes we can handle. This gives the above theorems, as well as some more technical (but more general) results in the body of the article, see Corollary 11.13, and Proposition 12.1. Instead of giving the full statements, we give some examples below to demonstrate them.
Figure 2: Some examples: unlabeled edges are assumed to have label 3.
@@
(=Corollary 12.4) All Artin groups whose Coxeter diagrams belong to Figure 2 satisfy the $K(\pi,1)$-conjecture.
All these are new for $K(\pi,1)$-conjecture. Figure 2 contains examples of any given dimension and examples with larger and larger irreducible spherical Artin subgroups. The main difference between examples here and some previous examples in [Hua24a, Figure 3] is the types of almost spherical sub-diagrams they are allowed to contain.
@@
Let $\Lambda$ be a connected Coxeter diagram satisfying the assumptions in one of the previous theorems. Then $A_{\Lambda}$ is torsion-free and it has a trivial center when $\Lambda$ is not spherical.
1.4 Non-positive curvature pieces in spherical Deligne complexes
Now we discuss the main ingredient in this article which is used to prove results on filling short cycles in Section 1.2, as well as some ideas in the proof. A standard situation when one knows how to fill in cycles in the space, and has strong control over the geometry and possibly combinatorics of the filling disk, is that the space satisfies a fine notion of non-positively curvature, like CAT$(0)$. So it would be ideal if all the spherical Deligne complexes for spherical Artin groups could be metrized as CAT$(0)$ spaces. However, this naive hope breaks down as spherical Deligne complexes are homotopic to wedges of spheres [Del72], so they can not admit CAT$(0)$ metric. Nevertheless, this does not rule out the possibility of finding large non-positively curved “islands” inside spherical Deligne complexes, and our strategy still works if some island is large enough to contain the cycle we are interested in. This situation bears some similarity to the geometrization of 3-manifolds/orbifolds - while a 3-manifold might not be metrized by a single Thurston geometry, it contains maximal “islands” which can be indeed metrized by a single Thurston geometry. This consideration leads us to the following question.
Question 1.11.
Given a spherical Artin group $A_{S}$ acting on its spherical Deligne complex $\Delta_{S}$. We want to find a subgroup $G\leq A_{S}$, and an $G$-invariant and $G$-cocompact subcomplex $Y\subset\Delta_{S}$ such that $Y$ admits an $G$-invariant metric of “non-positive curvature”. Moreover, we want $G$ and $Y$ to be as large as possible.
There is some flexibility of what we mean by “non-positive curvature” in Question 1.11. Besides CAT$(0)$, one may consider other forms of non-positive curvature [Lan13, DL15, CCHO14, CCG+20]. The bottom line is that such a form of non-positive curvature should give strong control on filling disks of cycles, so forms of non-positive curvature that are more “coarse” in nature might not work well for our purpose, see [Web20].
Now we describe the non-positively curved islands we found in detail. Suppose $A_{S}$ is irreducible and spherical. Let the associated pure Artin group be $PA_{S}$. Let $\overline{PA}_{S}$ be the central quotient of $PA_{S}$. As $A_{S}$ is spherical, we know $PA_{S}\cong\mathbb{Z}\times\overline{PA}_{S}$. Note that $\mathbb{Z}$ factor of $PA_{S}$ is canonical as it corresponds to the center of the group. However, the $\overline{PA}_{S}$ factor is not canonical - there are many different subgroups of $PA_{S}$ isomorphic to $\overline{PA}_{S}$, which are called $\overline{PA}_{S}$-sections in $PA_{S}$.
Recall that quotient $\Delta_{S}/PA_{S}$ is isomorphic to the Coxeter complex $\mathfrak{C}_{S}$. This gives a folding map $\pi:\Delta_{S}\to\mathfrak{C}_{S}$. Let $H$ be a wall in $\mathfrak{C}_{S}$ (i.e., fix points of a reflection), and let $K_{H}$ be the maximal subcomplex of $\mathfrak{C}_{S}$ contained in an open halfspace bounded by $H$. Then $\pi^{-1}(K_{H})$ is not connected, and the stabilizer of each connected component of $\pi^{-1}(K_{H})$ for the action $PA_{S}\curvearrowright\Delta_{S}$ is a $\overline{PA}_{S}$-section in $PA_{S}$.
A geodesic metric space is injective if whenever a collection of closed metric balls in the space have pairwise non-empty intersection, then their common intersection is non-empty. This is a form of non-positive curvature [Lan13].
Proposition 1.12.
(=Proposition 5.4) Suppose $A_{S}$ is of type $A_{n}$. Then for each wall $H$ of $\mathfrak{C}_{S}$, each component of $\pi^{-1}(K_{H})$ admits a piecewise $\ell^{\infty}$ injective metric that is invariant under the action of the associated $\overline{PA}_{S}$-section.
From a topological viewpoint, a component of $\pi^{-1}(K_{H})$ arises naturally from the deconing of the associated $A_{n}$-type arrangement of hyperplanes. When $\Delta_{S}$ is 2-dimensional, the component was considered by Falk in [Fal95] for a different purpose, with a different metric.  Though the description of the complex in Falk’s paper is different as well, it gives the same complex (see Section 4.3 for more details). Thus, from now on, we will refer to components of $\pi^{-1}(K_{H})$ as Falk subcomplexes of $\Delta_{S}$ associated with $H$.
There is a similar result about injective metric on Falk subcomplexes in type $B_{n}$ case, see Proposition A.1. We will not need Proposition A.1 in this article, although we will need it in another article, so we put it in the appendix.
Now we turn to the case that $A_{S}$ is of type $H_{3}$, which is the most challenging case in this article. Let $\pi:\Delta\to\mathfrak{C}_{S}$ and $H$ as before. Unfortunately, those Falk subcomplexes do not appear to admit forms of non-positive curvature (i.e. CAT$(0)$, injective, conformally CAT$(0)$, etc) which are invariant under the action of the associated sections. However, this is not fatal to our strategy. We will solve instead a variation of Question 1.11.
Question 1.13.
Given a spherical Artin group $A_{S}$ acting on its spherical Deligne complex $\Delta_{S}$. We want to find a subgroup $G\leq A_{S}$, and an $G$-invariant and $G$-cocompact subcomplex $Y\subset\Delta_{S}$ together with a quotient homomorphism $q:G\to\bar{G}$, a simplicial action of $\bar{G}$ on a simplicial complex $\bar{Y}$, and an $q$-equivariant quotient simplicial map $Y\to\bar{Y}$ such that $\bar{Y}$ admits an $\bar{G}$-invariant metric of “non-positive curvature”. Here we want $G$ and $Y$ to be as large as possible, and we want the fibers of $G\to\bar{G}$ and $Y\to\bar{Y}$ to be as small as possible.
The only difference between this question and Question 1.11 is that we do not require $Y$ to be non-positively curved on the nose. Instead, we are allowed to pass to an appropriate quotient $\bar{Y}$ of $Y$ which is non-positively curved. Answering this question still sheds lights on understanding minimal filling cycles in $\Delta_{S}$. Indeed, suppose we have a cycle $\omega$ that is in the “island” $Y$ of $\Delta_{S}$. Then we want to use the non-positive curvature in $\bar{Y}$ to fill in the image $\bar{\omega}$ of $\omega$, then try to lift at least part of the information of filling disk in $\bar{Y}$ back to $Y$. In order for the last step to work, we want the quotient map $Y\to\bar{Y}$ to lose as little information as possible.
When $A_{S}$ is of type $H_{3}$, we find several different pairs $\bar{G}$ and $\bar{Y}$, and combine their power to understand cycles in $\Delta_{S}$. The exact description of these pairs is a bit technical, which we refer to Section 6. Here we only explain the idea of finding such pairs. We take $G$ and $Y$ as before - $Y$ is a Falk subcomplex, and $G$ is the associated section. Here $G$ can be realized as the fundamental group of a manifold of form $\mathbb{C}^{2}-\cup_{\lambda\in\Lambda}H_{\lambda}$ where each $H_{\lambda}$ is an affine hyperplane in $\mathbb{C}^{2}$. If we consider a subset $\Lambda^{\prime}\subset\Lambda$, then the inclusion $\mathbb{C}^{2}-\cup_{\lambda\in\Lambda}H_{\lambda}\hookrightarrow\mathbb{C}^{2}-\cup_{\lambda\in\Lambda^{\prime}}H_{\lambda}$ naturally induces a quotient of $G$. The pairs $\bar{G}$ and $\bar{Y}$ are obtained by considering appropriate subsets $\Lambda^{\prime}$ of $\Lambda$.
It remains to ask, what if the 6-cycle $\omega$ in $\Delta_{S}$ is not contained in any “islands”? As these islands are chosen to be Falk subcomplexes, the failure of being contained in any islands implies the cycle has to be “spread-out” in an appropriate sense, hence can be handled effectively by the retraction map defined in [GP12, CP14, BP23, God23, DGM24] from an Artin group to its parabolic subgroups.
1.5 Structure of the article and reading guide
The article has three parts. Part 1 consists of Section 2, and it is a collection of preliminaries. Part 2 consists of Section 3 to Section 7, where we study 6-cycles in different types of spherical Artin complexes. More precisely, Section 3 is about the type $F_{4}$ case. Section 4 and Section 5 are about the type $A_{n}$ case. Section 6 and Section 7 are about the type $H_{3}$ case. Part 3 consists of Section 8 to Section 12, where we deduce new cases of the $K(\pi,1)$-conjecture.
Part 2 and Part 3 are mostly independent from each other. The main results of Part 2 are Proposition 3.2, Theorem 5.6, and Theorem 7.1. After reading Part 1, if the reader is willing to take these three main results of Part 2 for granted, then the reader can go to Part 3 directly, to see how these three results imply new cases of the $K(\pi,1)$-conjecture.
1.6 Acknowledgment
The author benefits greatly from several long and detailed conversations with the following colleagues on topics closely related to this article: Ruth Charney, Mike Davis, Thomas Haettel, Damian Osajda and Piotr Prytycki. The author also thanks Grigori Avramidi, Mladen Bestvina, Srivatsav Kunnawalkam Elayavalli, Michael Falk, Alexandre Martin, Jon McCommand, Katherine Goldman, Nima Hoda, Camille Horbez, Nick Salter, Rachel Skipper, Antoine Song, and Frank Wagner for helpful related conversation.
The author is partially supported by a Sloan fellowship and NSF DMS-2305411. The author thanks the Centre de Recherches Mathématiques in Montreal for the thematic program Geometric Group Theory in 2023 where part of the work was done. The author thanks the Institut Henri Poincaré (UAR 839 CNRS-Sorbonne Université) for the trimester program Groups acting on fractals, Hyperbolicity and Self-similarity in 2022 (through LabEx CARMIN, ANR-10-LABX-59-01) where part of the work was done.
2 $K(\pi,1)$-conjecture via relative Artin complexes
This is a dual-purpose section. It serves as a summary of preliminaries for the whole article, as well as a review of an approach to $K(\pi,1)$-conjecture in [Hua24a] which provides motivation for the preliminary material. The presentation of the approach to $K(\pi,1)$-conjecture here has a few new parts compared to [Hua24a]. While some parts of the strategy are still under construction, the main frame is already transparent, and it is simple.
2.1 Artin complexes and relative Artin complexes
Let $W_{S}$ and $A_{S}$ be the Coxeter group and Artin group with generating set $S$. If $\Lambda$ is a Coxeter diagram, then we will also write $W_{\Lambda}$ and $A_{\Lambda}$ for the Coxeter group and Artin group with Coxeter diagram $\Lambda$. There is a homomorphism $A_{S}\to W_{S}$, whose kernel is called the pure Artin group, and is denoted by $PA_{S}$.
Any $S^{\prime}\subset S$ generates a subgroup of $A_{S}$ which is also an Artin group, whose Coxeter diagram is the full subgraph $\Lambda_{S^{\prime}}$ of $\Lambda$ spanned by $S^{\prime}$ [vdL83]. This subgroup is called a standard parabolic subgroup of type $S^{\prime}$. A parabolic subgroup of $A_{S}$ of type $S^{\prime}$ is a conjugate of a standard parabolic subgroup of type $S^{\prime}$. A parabolic subgroup of $A_{S}$ is reducible if its type $S^{\prime}$ admits a disjoint non-trivial decomposition $S^{\prime}_{1}\sqcup S^{\prime}_{2}$ such that each element in $S^{\prime}_{1}$ commutes with every element in $S^{\prime}_{2}$. If such decomposition does not exist, then the parabolic subgroup is irreducible.
Recalled that the Artin complex, introduced in [CD95] and further studied in [GP12, CMV20], defined as follows. For each $s\in S$, let $A_{\hat{s}}$ be the standard parabolic subgroup generated by $S\setminus\{s\}$. Let $\Delta_{S}$ be the simplicial complex whose vertex set is corresponding to left cosets of $\{A_{\hat{s}}\}_{s\in S}$. Moreover, a collection of vertices span a simplex if the associated cosets have a nonempty common intersection. It follows from [GP12, Proposition 4.5] that $\Delta_{S}$ is a flag complex. The Artin complex is an analog of Coxeter complex in the setting of Artin group. The definition of a Coxeter complex $\mathfrak{C}_{S}$ of a Coxeter group $W_{S}$ is almost identical to Artin complex, except one replaces $A_{\hat{s}}$ by $W_{\hat{s}}$, which is the standard parabolic subgroup of $W_{S}$ generated by $S\setminus\{s\}$. Each vertex of $\mathfrak{C}_{S}$ or $\Delta_{S}$ corresponding a left coset of $W_{\hat{s}}$ or $A_{\hat{s}}$ has a type, which is defined to be $\hat{s}=S\setminus\{s\}$. The type of each face of $\mathfrak{C}_{S}$ or $\Delta_{S}$ is defined to be the subset of $S$ which is the intersection of the types of the vertices of the face. In particular, the type of each top-dimensional simplex is the empty set. The quotient complex of $\Delta_{S}$ under the action of the pure Artin group is isomorphic to $\mathfrak{C}_{S}$.
Note that if each $A_{\hat{s}}$ is spherical, but $A_{S}$ is not spherical, then $\mathbb{D}_{S}$ is isomorphic to the barycentric subdivision of $\Delta_{S}$. When $A_{S}$ is spherical, then sometimes $\Delta_{S}$ is also called spherical Deligne complex - indeed, let $\mathcal{A}$ be the reflection arrangement associated with the spherical Artin group $A_{S}$. Then the spherical Deligne complex $\mathbb{S}_{\mathcal{A}}$ defined in Section 4.3 is isomorphic to $\Delta_{S}$.
Definition 2.1.
([Hua24a]) Let $A_{S}$ be an Artin group with Coxeter diagram $\Lambda$. Let $S^{\prime}\subset S$. The $(S,S^{\prime})$-relative Artin complex $\Delta_{S,S^{\prime}}$ is defined to be the induced subcomplex of the Artin complex $\Delta_{S}$ of $A_{S}$ spanned by vertices of type $\hat{s}$ with $s\in S^{\prime}$. In other words, vertices of $\Delta_{S,S^{\prime}}$ correspond to left cosets of $\{A_{\hat{s}}\}_{s\in S^{\prime}}$, and a collection of vertices span a simplex if the associated cosets have nonempty common intersection.
Let $\Lambda^{\prime}$ be the induced subgraph of $\Lambda$ spanned by $S^{\prime}$. Then we will also refer an $(S,S^{\prime})$-relative Artin complex as $(\Lambda,\Lambda^{\prime})$-relative Artin complex, and denote it by $\Delta_{\Lambda,\Lambda^{\prime}}$.
The links of vertices in relative Artin complexes can be computed via the following result from [Hua24a, Lemma 6.4].
Lemma 2.2.
Let $\Delta$ be the $(\Lambda,\Lambda^{\prime})$-relative Artin complex, and let $v\in\Delta$ be a vertex of type $\hat{s}$ with $s\in\Lambda^{\prime}$. Let $\Lambda_{s}$ and $\Lambda^{\prime}_{s}$ be the induced subgraph of $\Lambda$ and $\Lambda^{\prime}$ respectively spanned all the nodes which are not $s$. Then the following are true.
  1. 1.
  2. 2.
  3. 3.
@@ 2.1 Artin complexes and relative Artin complexes
([Hua24a, Corollary 6.5]) Let $\Delta$ be the $(\Lambda,\Lambda^{\prime})$-relative Artin complex. For $i=1,2,3$, let $x_{i}\in\Delta$ be a vertex of type $\hat{s}_{i}$ with node $s_{i}\in\Lambda^{\prime}$. Suppose $s_{1}$ and $s_{3}$ are in different components of $\Lambda\setminus\{s_{2}\}$.
If $x_{1}$ is adjacent to $x_{2}$, and $x_{2}$ is adjacent to $x_{3}$, then $x_{1}$ is adjacent to $x_{3}$.
Lemma 2.4.
([Hua24a, Lemma 7.1]) Let $\Lambda_{1}\subset\Lambda_{2}$ be two induced subgraphs of $\Lambda$ such that $\Lambda_{2}\setminus\Lambda_{1}$ contain exactly one node, denoted by $s$. If $\operatorname{lk}(x,\Delta_{\Lambda,\Lambda_{2}})$ is contractible for some (hence any) vertex $x\in\Delta_{\Lambda,\Lambda_{2}}$ of type $\hat{s}$, then $\Delta_{\Lambda,\Lambda_{2}}$ deformation retracts onto $\Delta_{\Lambda,\Lambda_{1}}$.
2.2 $K(\pi,1)$-conjecture via relative Artin complexes
Our point of departure is the following.